OpenStudy (anonymous):

the ratio of fanatics to zealots was 11 to 4. if there were 2280 fanatics and zealots in all, how many were fanatics?

OpenStudy (anonymous):

11x = 4y x +y = 2280 11/4x = y 15/4x = 2280 x = 608, y = 1672

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

teams ratio of games won to games played was 2 to 5. If the team played 55 games, how many games did the team lose?

OpenStudy (anonymous):